OpenStudy (llort):

Today I cycled up a steep slope, the slope started at the 1.6 mile mark of my journey and ended at the 3.4 mile mark. The elevations ranged from 781 ft above sea level to 1086 ft. How steep was the slope?

OpenStudy (matt101):

Assuming a constant slope (otherwise we're taking the average slope), it's the change in rise over the change in run: (1086-781)/(3.4-1.6) =295/1.8 =163.8 That's one steep slope :)

OpenStudy (llort):

does it matter that one is in miles and the other is in feet?

OpenStudy (dumbcow):

use slope formula m = (1086 - 781)/(3.4-1.6) = 305/1.8 = 169.44 ft/mi or divide by 5280 = .0321

OpenStudy (matt101):

Actually I didn't consider that...probably, then the slope would be a more reasonable number.

OpenStudy (llort):

I have a feeling that this is a trig question, treat the distance traveled as a hypotenuse and the elevation as the opposite side.

OpenStudy (radar):

me too.

OpenStudy (radar):

slant range!

OpenStudy (matt101):

Could be, but they would have to give you some way of finding an angle, which they don't

OpenStudy (llort):

I am just giving you all the data from my GPS

OpenStudy (dumbcow):

then sinx = 305/9504 x = sin^-1(305/9504) x = 1.839 degrees

OpenStudy (matt101):

Oh lol real life situation :)

OpenStudy (radar):

was the 1.6 and 3,4 mark from GPS or your cycle odometer?

OpenStudy (llort):

my GPS, it is correct within 4 yards

OpenStudy (radar):

O.K just curious, the odometer would of been slant range. don't know what the GPS is measuring: actual range or displacement??

OpenStudy (llort):

yeah, that could change the answer

OpenStudy (radar):

probably delta X

OpenStudy (dumbcow):

actually it doesn't change the degree of elevation by much if it is only measuring displacement tan(x) = 305/9504 x = 1.838 degrees

OpenStudy (llort):

the slope looked much steeper... strange

OpenStudy (dumbcow):

thats almost a 2% grade so yeah pretty steep for cycling good for you :)

OpenStudy (llort):

thanks :) it registered on mapmyride as a category 5 climb.